OpenStudy (anonymous):

Triangle ABC is a right triangle. If AB = 3 and AC= 7, find BC Leave answer in simplest radical form

OpenStudy (caitlinnr14):

what do you think you should do first

jhonyy9 (jhonyy9):

can you drawing it ?

OpenStudy (caozeyuan):

let A be the right angle, solve for BC, then let B be the right angle, solve for BC a second time, now let C be the right angle, solve for BC the third time. So you have 3 different values of BC cuz any angle can be the right angle

jhonyy9 (jhonyy9):



jhonyy9 (jhonyy9):

so if AB = 3 and AC = 7 how you find BC ?

jhonyy9 (jhonyy9):

do you know the Pythagoras theorem in case of a right triangle what said ?

jhonyy9 (jhonyy9):

@caitlinnr14 can you solve it for medal ?

jhonyy9 (jhonyy9):

so using this theorem you can writing AC^2 = AB^2 + BC^2 from this you need calculing BC than BC^2 = AC^2 - AB^2 BC^2 = 49 -9 BC^2 = 40 BC = sqrt(40) BC = sqrt(4*10) BC = 2sqrt10